Write equations showing how each of the following weak bases ionizes water to form OH−.

1) NH3
2) HCO−3
3) CH3NH2

Consider a 0.10 M solution of a weak polyprotic acid (H2A) with the possible values of Ka1 and Ka2 given below. Calculate the contributions to [H3O+] from each ionization step.

1)Ka1=1.0×10^−4;Ka2=5.0×10^−5
2)Ka1=1.0×10^−4;Ka2=1.0×10^−5
3)Ka1=1.0×10^−4;Ka2=1.0×10^−6
1
Expert's answer
2020-04-06T12:06:11-0400

1)Ktotal= 1.0×10^−4 x 5.0×10^−5 = 5e-9

2)Ktotal=1.0×10^−4 x 1.0×10^−5 = 1e-9

3)Ktotal=1.0×10^−4 x 1.0×10^−6=1e-10